13. Mr Fu asked a garden designer to make a square flower bed. The cost for paving the border of the flower bed is $20 per metre. The cost of the flowers and soil in the flower bed is $60 per metre^2. If the total cost of making the flower bed is $3500, find the length of a side of the flower bed.
Answer by ankor@dixie-net.com(22740)

Mr Fu asked a garden designer to make a square flower bed.
The cost for paving the border of the flower bed is $20 per metre.
The cost of the flowers and soil in the flower bed is $60 per metre^2.
If the total cost of making the flower bed is $3500, find the length of a side of the flower bed.
:
Let x = length of the side of the flower bed
then
60x^2 = cost of the flower bed
and
20(4x) = 80x = cost of the border
:
Bed cost + border cost = total cost
60x^2 + 80x = 3500
A quadratic equation
60x^2 + 80x - 3500 = 0
Simplify divide by 20, results:
3x^2 + 4x - 175 = 0
Factor this to
(3x + 25)(x - 7) = 0
positive solution
x = 7 meters side of the garden
:
:
Check solution by finding the cost
60(7^2) = 2940
20(4*7) = 560
----------------
total$: 3500